6 children earned 42-five peso coins and 24- ten peso coins from selling newspaper. if they share the coins equally, what fraction of the total coins should each get?
Found 2 solutions by CubeyThePenguin, josgarithmetic:
Answer by CubeyThePenguin(3113)

total number of coins: 42 + 24 = 66
Each child gets 6/66 = 1/11 of the total.

Answer by josgarithmetic(39799)

Both of the coin counts are integer multiples of 6.